OpenStudy (elleblythe):

As a function of theta where 0

OpenStudy (mathmath333):

\(\Large \rm cos\theta=(360+\theta)\)

OpenStudy (mathmath333):

\(\Large\rm 427=360+67 \)

OpenStudy (mathmath333):

\(\Large\rm cos67=0.39 \)